A LED company claims that the average life of the LED light bulbs it

manufactures is 1, 500 hours with a standard deviation of 500 hours. If a random sample of 40 bulbs is chosen, what is the probability that the sample mean will be:


a. greater than 1, 400 hours?


b. less than 1, 400 hours?


1
Expert's answer
2021-05-31T00:04:56-0400

Let "X=" the average life of the LED light bulb: "X\sim N(\mu, \sigma ^2/n)."

Given "\mu=1500\ h, \sigma=500\ h, n=40"


a)

"P(X>1400)=1-P(\leq1400)"

"=1-P(\leq\dfrac{1400-1500}{500/\sqrt{40}})\approx1-P(Z\leq-1.264911)"

"\approx0.897048"

b)

"P(X<1400)=P(Z<\dfrac{1400-1500}{500/\sqrt{40}})"

"\approx P(Z<-1.264911)\approx0.102952"